what is 9.21 g of barium in moles?
1
Expert's answer
2018-01-16T01:58:53-0500
Solution:
The molar mass of Ba is 137.7 g/mol
n(Ba) = m(Ba) / Mr(Ba)
n(Ba) = 9.21 / 137.7 = 0.067 mol
Answer:
n(Ba) = 0.067 mo